DEPRECIATION METHODS

Charlene is evaluating a capital budgeting project that should last for 4 years. The project requires $975,000 of equipment. She is unsure what depreciation method to use in her analysis, straight-line or the 3-year MACRS accelerated method. Under straight-line depreciation, the cost of the equipment would be depreciated evenly over its 4-year life (ignore the half-year convention for the straight-line method). The applicable MACRS depreciation rates are 33%, 45%, 15%, and 7%. The company's WACC is 13%, and its tax rate is 35%.

  1. What would the depreciation expense be each year under each method? Round your answers to the nearest cent.
    Year Scenario 1
    (Straight-Line)
    Scenario 2
    (MACRS)
    1 $ $
    2
    3
    4
  2. Which depreciation method would produce the higher NPV?
    Straight-Line OR MACRS?

    How much higher would the NPV be under the preferred method? Round your answer to two decimal places. Do not round your intermediate calculations.
Answer #1

Depreciation is the reduction in the value of an asset due to its wear and tear. Depreciation can be calculated using two methods-

a) Straight Line method- it is the most simplest method of calculating depreciation. In this method, the value of the asset is reduced uniformly over each period until it reaches its salvage value. The value of depreciation remains the same.

b) MACRS Depreciation- The modified accelerated cost recovery system (MACRS) is a depreciation system used for tax purposes. MACRS allows for faster depreciation in the first years of an asset's life and slows depreciation later on. This is beneficial to businesses from a tax perspective.

In the above question,

Cost of asset= $975,000 Life of asset,n= 4 years

WACC-13% tax rate,t=35%

A)

Years Cost of asset Dep rate(MACRS) Straight line ($) MACRS ($)
1 975000 0.33 975000/4=243750 975000*0.33=321750
2 975000 0.45 975000/4=243750 975000*0.45=438750
3 975000 0.15 975000/4=243750 975000*0.15=146250
4 975000 0.07 975000/4=243750 975000*.07=68250

B) NPV using Straight line method:

WACC= 0.13   

YEAR DEPRECIATION VALUE TAX CASH FLOW PV
1 243750 0.35 243750*0.35=85312.5 =85312.5/(1+0.13)^1=75497.59
2 243750 0.35 243750*0.35=85312.5 =85312.5/(1+0.13)^2=66812.2
3 243750 0.35 243750*0.35=85312.5 =85312.5/(1+0.13)^3=59125.84
4 243750 0.35 243750*0.35=85312.5 =85312.5/(1+0.13)^4=52323.75

Sum of PV of cash inflows= 75497.59+ 66812.2+59125.84+52323.75 = 253759.6

NPV= PV of cash inflows- PV of cash outflows

= 253759.6-975000 = -721240.41

NPV using MACRS

WACC= 0.13   

YEAR DEPRECIATION VALUE TAX CASH FLOW PV
1 321750 0.35 321750*0.35=112612.5 =112612.5/(1+0.13)^1=99657.08
2 438750 0.35 438750*0.35=153562.5 =153562.5/(1+0.13)^2=120621.96
3 146250 0.35 146250*0.35=51187.5 =51187.5/(1+0.13)^3= 35475.50
4 68250 0.35 68250*0.35=23887.5 =23887.5/(1+0.13)^4= 14650.65

Sum of PV of cash inflows= 99657.08+ 120621.96+35475.50+14650.65 = 270045.19

NPV= PV of cash inflows- PV of cash outflows

= 270045.19-975000 = -704954.80

The NPV produced from MACRS will be higher by ( 721240.41- 704954.80) = 16285.61

NOTE :

Depreciation is a tax deductible expense. With depreciation, we save on paying income tax. So, the tax saving is a cash inflow.

If Depreciation is not tax deductible, then we would have to pay higher income tax.